# Semantic Structure-Mapping in LLM and Human Analogical Reasoning

### v1

@@ 3.1.1 Overall Performance
Table 1: Defaults and control conditions used to measure ability of humans and LLMs to perform analogical reasoning tasks that involve semantic structure-mapping. The Permuted Questions condition (not shown) is identical to Defaults, but with question order permuted.
@@
Figure 2 shows the performance of humans and LLMs in the Defaults condition as a function of their performance on MMLU [^2] MMLU scores are few-shot for GPT-4 and 5-shot for other models. The reported human baseline is the estimate for human experts given by Hendrycks et al. [28]. The score for Pythia 12B could not be found and so we use the reported value for Pythia 6.9B Tulu., a widely-used language competency benchmark. Increasing MMLU score is associated with higher accuracy on the Defaults condition. Smaller models do not perform competitively (Pythia-12B obtains an accuracy of 0.0, Falcon 40B 0.1, GPT-3 0.5, and Claude 2 0.6). This steadily increasing performance is presumed to correlate with the scale of model parameters and training data [29]. We focus our remaining analysis on comparing human subjects to GPT-4 and Claude 3. In the Defaults condition, neither GPT-4 (coef=-0.7696, z=-1.659, p=0.097) nor Claude 3 (coef=-0.6131, z=-1.299, p=0.194) performs significantly worse than human subjects.
@@
Figure 2: Human and LLM accuracy in the Defaults condition, relative to performance on the MMLU benchmark. Models in blue are not instruction-tuned while models in orange are. Error bars show standard errors.
@@
Figure 3 compares humans to high-performing LLMs in the Defaults and Permuted Pairs conditions. LLM performance drops in the Permuted Pairs condition, while humans seem equally able to infer the mapping regardless of word presentation order. This effect is significant for both Claude 3 (coef = -1.7802, z = -4.217, p < 0.001) and GPT-4 (coef = -1.6796, z = -3.975, p < 0.001). This suggests that, while the overall performance is comparable, there are likely meaningful mechanistic differences in how the analogy is processed in humans versus LLMs. The remaining control conditions and data for all tested models are shown in Figure 15 of the Appendix . In these conditions, we find that humans and models are roughly equally affected. For example, accuracy in the Distracted condition drops by approximately 0.25 for all three subject types.
@@
Figure 3: Human and LLM accuracy in the Defaults and Permuted Pairs conditions. Error bars show standard errors.
@@ 3.1.2 Effect of Semantic Structure on Reasoning
We next investigate more directly the extent to which humans and LLMs leverage semantic structure in order to complete our analogy tasks. To do this, we design three variants of our Defaults analogy task (see Table 2). First, the Only RHS condition removes the source domain entirely. High performance in this condition thus indicates that a subject is able to complete the questions based only on the evident pattern in the target domain. We then introduce two variants which make the semantic structure in the source domain less coherent: the Randoms condition uses unrelated words, while the Random Finals condition uses of three related words followed by one random word. We thus take the performance difference between the RHS Only condition and either the Random or Random Final condition to be a measure of the subject’s bias toward using the semantic structure of the source domain. That is, if the subject is capable of solving the task by simply ignoring the left hand side (the Only RHS condition), then poor performance in the other conditions indicates that the subject was misled by the presence of the altered left hand side.
Only RHS Test of how well the answer can be inferred without using any structure-mapping C C C c c c C C
Random Last Variant of Defaults in which the final term is not semantically related to the preceding terms square => C C C rectangle => c c c circle => C C lime =>
@@
Both humans and models competently complete the Only RHS condition (see Figure 4). Accuracy is approximately 0.8 for Claude 3 with human subjects and GPT-4 slightly higher at 0.9. GPT-4 is not significantly different from humans in this condition (coef = 0.1178, z = 0.223, p = 0.824), and Claude 3 is worse than humans by a barely significant margin (coef = -0.9130, z = -1.994, p = 0.046). Thus, both humans and LLMs are able to complete the task without the guidance of the left hand side. Considering this, we look at the performance degradation associated with encountering incoherent semantic structure on the left hand side. Humans exhibit a modest decrease in accuracy of about 0.15 in the Random and Random Permuted Pairs conditions relative to defaults. Claude-3 and GPT-4, however, exhibit much larger drops: Claude 3 decreases by approximately 0.5 relative to Defaults, while GPT-4 decreases by 0.6 and 0.4 in the Random and Random Permuted Pairs conditions. Across these two conditions, both GPT-4 (coef = -2.1972, z = -5.211, p < 0.001) and Claude 3 (coef = -2.0680, z = -4.960, p < 0.001) perform significantly worse than humans.
@@
Figure 4: Human and LLM accuracy in Only RHS, Randoms, and Random finals conditions. Data from the Random Permuted Pairs condition is shown in Figure 15 of the Appendix . Error bars show standard errors.
@@
From this we conclude that human subjects are able to easily identify when the left hand side contains no useful semantic structure to leverage. When there is none, they are able to employ a strategy that only relies on the right hand side. By contrast, models do not seem capable of easily identifying the lack of informativeness of the left hand side in these conditions, as they do not use the strategy of only attending to the right hand side, even though they show their capability of using this strategy when no left hand side is present. This suggests mechanistic differences between how human subjects and models process this task.
@@
Although the performance of human subjects does not drop notably in the Random condition compared to the Only RHS condition, it does drop by a wide margin in the Random Finals condition. In this condition, accuracy is approximately 0.5 lower than in the Only RHS condition. This further suggests that the semantic relatedness of the left hand side affects the strategy of human subjects: when the left hand side is clearly unrelated, the information it provides is discarded, but when much of the left hand side appears related, the information is not discarded and the random final word of the source domain prompts an incorrect answer from human subjects. Models also show a large drop in performance in the Random Finals condition relative to Only RHS, with Claude 3 dropping by 0.5 and GPT-4 dropping by 0.8. Simple effects analysis shows that both Claude 3 (coef = -1.0464, z = -2.799, p = 0.005) and GPT-4 (coef = -2.7850, z = -5.168, p < 0.001) are significantly worse than humans in the Random Finals condition. However, we see this difference as less informative than that both models drop in performance across all the random conditions relative to their own performance in the Only RHS condition.

### v2

@@ 2.1 Study 1: Analogical Inference from Semantic Structure
Table 1: Defaults and control conditions used to measure ability of humans and LLMs to perform analogical reasoning tasks that involve analogical reasoning from a linguistic to a non-linguistic domain. The Permuted Questions condition (not shown) is identical to Defaults, but with question order permuted.
Our second set of controls is designed to diagnose the extent to which subjects actually leverage semantic structure in the source domain to solve these tasks.
The Only RHS condition removes the words entirely, presenting only the symbol sequences. This condition tests whether subjects can complete the patterns through purely formal reasoning without semantic guidance.
The Randoms condition uses semantically unrelated words in the source domain, testing whether subjects will identify the semantic content on the left hand side as “misleading” and thus focus exclusively on symbolic patterns in the target domain.
By contrast, the Random Finals condition initially establishes a semantic pattern but breaks it in the final term, testing whether subjects will dynamically shift their strategy when semantic structure appears to be “misleading.”
We interpret a performance discrepancy between either the Random or Random Finals conditions and the Only RHS condition as evidence that the subject has a bias toward using semantic structure even when ignoring it altogether would enable a simpler representation of the pattern. See Table 2 for examples from these conditions.
Only RHS Test of how well the answer can be inferred without using the source domain C C C c c c C C
Random
Finals
Variant of Defaults in which the final term is not semantically related to the preceding terms square => C C C rectangle => c c c circle => C C lime =>
The $2\times n$ condition provides the strongest test of schema induction by requiring the integration of multiple relationships to determine the correct mapping (see the example in Table 3). For instance, subjects might need to understand both what body parts are covered by what clothing items in the source domain and how this relationship determines symbol patterns in the target domain. This condition prevents a solution through simple pattern completion or one-to-one mapping, instead requiring subjects to induce and apply a higher-order relational rule.
$2\times n$ A condition in which words from a variety of consistent pairs are shown, with a single complete pair (‘torso’ and ‘shirt’ in the example) revealing how the relation is encoded on the RHS. pants => H # H glove => X # X torso => V foot => Z head => M shirt => V # V hat =>
Table 3: The $2\times n$ condition, used to diagnose subjects’ tendency to rely on RHS-only heuristics to solve the task. In this example, reasonable answers based on only the right hand side include H, X, Z#Z, and M#M to complete pairs in accordance with the relationship between V and V#V. These four responses give a baseline RHS-only success of 25%. All questions in this condition share this feature of having a low baseline probability of success if only the right hand side is used.
2.2 Study 2: Analogical Inference from Semantic Content
While Study 1 focused on mapping structural relations, Study 2 investigates how subjects identify relevant semantic properties of individual concepts and induce rules that systematically connect these properties to formal patterns. Importantly, none of the tasks of Study 2 can be solved through pure pattern matching by considering only symbol sequences in the target domain.
The study includes four conditions that vary the type and number of semantic properties governing the mappings. Each condition requires subjects to identify relevant semantic properties and use them to guide inference, rather than applying transformation rules as in Study 1.222While participants could use a property-based approach to solve the Semantic Structure problems, the Semantic Content problems are differentiated in that this is the only approach that will yield the reference answers.
In the Categorial condition, subjects must identify a single categorial distinction that determines which symbol is used in the mapping. For example, participants might see various animals mapped to either ‘*’ or ‘!’, where the pattern depends on whether each animal is a mammal. Success requires searching through possible categorial distinctions (such as vertebrate/invertebrate, flying/non-flying, or mammal/non-mammal) to identify which one accounts for the observed symbol assignments.
The Dual-attribute condition increases complexity by requiring participants to track two independent categorial properties simultaneously. These properties jointly determine both the type of symbol(s) used in the target domain and the pattern they form. For example, subjects might see family members mapped to sequences of symbols where gender determines the symbol type (male $\rightarrow$ ‘!’, female $\rightarrow$ ‘*’) and generation number determines the sequence length (grandparent $\rightarrow$ 1, parent $\rightarrow$ 2, child $\rightarrow$ 3). This condition tests participants’ ability to decompose complex patterns into independent components and identify relevant semantic features for each component.
The Numeric condition focuses specifically on quantitative semantic properties, requiring participants to identify numeric features that determine the length of symbol sequences. For example, participants might see various animals mapped to sequences of asterisks where the sequence length matches the number of legs each animal has. This condition tests subjects’ abilities to re-represent concepts in the source domain in terms of their quantitative properties and identify which specific quantitative property explains the observed formal patterns in the target domain.
Finally, the Numeric Dual-attribute condition combines categorial and quantitative reasoning. Subjects must simultaneously identify a categorial feature that determines symbol choice and a numeric feature that determines sequence length. Like the Dual-attribute condition, this condition is particularly demanding as it requires inducing and integrating two independent rules while filtering out irrelevant properties of the concepts. These conditions are shown in Table 4.
Categorial Right-hand terms are single characters corresponding to a Categorial property of the left-hand terms. chicken => ! spider => ! cat => * horse => * ant => ! dog => * bee => ! human => Dual-attribute Right-hand terms are a sequence of several characters that vary according to two properties of the left-hand terms. grandfather => ! grandmother => * mother => * * father => ! ! brother => ! ! ! sister => Numeric Right-hand terms are a sequence of a single repeated character, with the number of repetitions corresponding to a numeric property of the left-hand terms. chicken => * * human => * * dog => * * * * spider => * * * * * * * * cat => * * * * horse => * * * * bee => Numeric
Dual-attribute
Right-hand terms are a sequence of a repeated character, with the number of repetitions corresponding to a numeric property of the left-hand terms and the character corresponding to a Categorial property. horse => * * * * cat => * * * * ant => ! ! ! ! ! ! bee => ! ! ! ! ! ! chicken => ! ! spider => ! ! ! ! ! ! ! ! dog => * * * * human =>
Table 4: The conditions of the Semantic Content experiment.
Across all conditions, the central challenge is again that of flexible re-representation: participants must dynamically adjust how they represent familiar concepts in the source domain based on the patterns they observe in the target domain. This requires suppressing typical ways of thinking about concepts in favor of task-relevant features. For example, when thinking about animals, one might typically focus on habitat, diet, or behavior, but the task might require representing them solely in terms of anatomical features like number of legs or reproductive characteristics.
3 Methods
3.1 Experiment Details
In the Semantic Structure experiment, each subject was presented with a quiz, which is a sequence of four such questions generated using four sets of base domains and four sets of target domains selected such that a participant sees each base and target domain exactly once. Eight variants of the task were devised to investigate the influence of task variations as described above.
Questions are introduced with the prompt “We are conducting an experiment on general reasoning abilities. Below we will show you various words and drawings of each, after which you will need to complete the last drawing. Respond as concisely as possible with only the last drawing.” Here, the term “drawings” directs the subject to complete the right hand side on the basis of the left, without directly referencing “analogical reasoning”. The experiment is introduced to human subjects and LLMs as studying “general reasoning abilities” for the same reason.
In the Semantic Content experiment, each condition (described in Table 4) contains two quizzes, with four questions per quiz. Unless otherwise stated, methodological details of the Semantic Content experiment match those of the Semantic Structure experiment.
3.2 Participants
3.2.1 LLMs
We run our experiments on the following LLMs: GPT-3 (text-davinci-002) [20], GPT-4 (gpt-4-0314) [21], Pythia-12B [22], Claude 2 [23], Claude 3 Opus [24], Falcon-40B [25], and Llama-405B [26]. All of the above are transformer-based LLMs trained primarily on a next-word prediction objective. OpenAI models are queried through the OpenAI API, Anthropic models are queried through the Anthropic API, Pythia is run on a local GPU server, and Llama is run through the Fireworks API. Across all models, we use a temperature of 0.7, and top_p is set to 1.
GPT-3 consists of a 175B parameter model trained on text completion and fine-tuned to produce more coherent answers. The details of GPT-4 are not publicly known, but it is considered by some sources to be a mixture-of-experts model consisting of numerous GPT-3-scale language models [27]. GPT-4, unlike the GPT-3 Davinci-002 model used for testing, supplements text-completion pre-training and fine-tuning with reinforcement learning from human feedback (RLHF) in order to better align model outputs with the expectations of a human user. The Davinci-002 version of GPT-3 employs a “similar but slightly different” response alignment method to RLHF, also using human data [28, 29]. The training of Claude 2 includes RLHF, but its performance falls short of GPT-4. The more recent Claude 3 (in our case, the most advanced Opus version) is considered to approximately match GPT-4 performance in general. GPT-3 and -4 are developed by OpenAI, whereas Claude 2 and 3 are developed by Anthropic. Pythia-12B and Falcon-40B are open-weights LLMs trained on a text-completion objective and consist of 12B and 40B parameters respectively. Neither undergoes RLHF. Pythia-12B is developed by EleutherAI, and Falcon-40B is developed by the Technology Innovation Institute. The 405B parameter Llama 3 model, developed by Meta, is among the best available open-weights models. It is a relatively straightforward transformer architecture and is not a mixture of experts. It undergoes post-training based on supervised fine-tuning and other techniques.
3.2.2 Human Subjects
We also test human participants on our experiments. Reported in the main text are results obtained from 194 (mostly undergraduate) Brown University students (132 in the Semantic Structure experiment, and 62 in the Semantic Content experiment). The split of participants between experiments approximately matches the 9:4 ratio of experiment conditions. The number of participants by condition are as follows: Defaults 18, Distracted 18, Only RHS 18, Permuted Pairs 17, Permuted Questions 17, Random Finals 15, Random Permuted Pairs 6, Randoms 8, $2\times n$ 15, Categorial 16, Dual-attribute 16, Numeric 16, Numeric Dual-attribute 14. The Defaults, Distracted, Only RHS, Permuted Pairs, Permuted Questions, and Random Finals conditions each have four quizzes per condition while the remaining conditions each have two quizzes. Subjects were assigned randomly to a single quiz from one condition without the re-use of subjects. Roughly the same number of participants were assigned to each condition, with the exception of the Randoms and Random Permuted Pairs conditions. These were together assigned roughly the expected number of subjects for a single condition due to their similarity.
3.3 Statistical testing
In each experiment, we are interested in the relative performance of human subjects and the best-performing models and how this depends on the particular experiment conditions. Differences between most models and human subjects are large and do not require statistical analysis, and so we focus our statistical analysis on the performance of GPT-4, Claude 3, and Llama-405B each relative to human subjects.
For each experiment and pair of subjects (human subjects and GPT-4, human subjects and Claude 3) we fit a logistic model to the data with and without interactions between the subject type and the experiment condition. In all cases, the outcome variable is the un-aggregated per-question score achieved by a subject (either a 0 or 1), and the predictor variables are experiment condition (e.g. “Defaults” or “Permuted Pairs”) and subject type (e.g. “human subjects” or “GPT-4”). We use six likelihood ratio tests to assess whether the interaction between subject type and experiment condition is significant for a given pair of subjects within a particular experiment, as motivated by Glover [30]. In all cases, the interaction is significant, and so we use simple effects analysis to investigate the direction and significance of the effect of subject type within particular conditions.
For the Semantic Content experiment, we additionally perform a logistic simple effects analysis comparing the performance of a single subject type (human, GPT-4, Claude 3, or Llama-405B) in compositional versus non-compositional conditions for the numeric and non-numeric cases respectively with the non-compositional condition as reference. For example, we assess the effect of the condition being Dual-attribute with Categorial as the reference condition for only the subject type Claude 3 (and likewise for the other three examined subject types).
Standard errors are computed first per subject type per question number and then propagated to the condition level for each subject type. To find the standard error per question, we aggregate over the responses of different individuals; for models we aggregate over multiple samples from the same model.
Further details are provided in Sections A.1 and A.2 of the Appendix.
4 Results
4.1 Reasoning From Semantic Structure
Recall (from §2.1) that our Semantic Structure experiments are designed to assess analogical inference by asking subjects to map a set of words containing relational structure in the source domain to a set of strings related via non-linguistic string operations in the target domain. We describe both humans’ and LLMs’ overall performance first (§4.1.1) before interpreting how performance changes within each of the more nuanced control conditions (§4.1.2–§4.1.5).
4.1.1 Overall Performance
Human subjects perform well overall in the Defaults setting, with a proportion of between 0.4 and 0.9 of answers matching the reference across the various conditions. [^3] We use “match to reference” instead of “accuracy” as there is, by definition, room for interpretation within analogy making and therefore no objectively correct solutions. For analysis purposes, we grade responses against specific reference answers. We note that, in principle, since subjects receive 4 analogy questions per quiz, the reference solution should be preferred over other justifiable solutions after seeing multiple examples. The most advanced LLMs that we test match references in the range 0.1-0.95 in the different conditions we assess. This performance range is comparable to prior work on analogical reasoning over arbitrary symbols. For example, the results of human subjects on the “zero-generalization setting” studied by both Webb et al. [7] and Lewis & Mitchell [9] range from 0.2-0.8 in the former study and from 0.5-1.0 in the latter study. Similarly, results for LLMs (GPT-3, GPT-3.5, and GPT-4) across those conditions range from 0.1-1.0 in the two studies. Thus, our data suggest that our analogy problems are not inherently easier or harder than those used in prior work.
@@ 4.1.1 Overall Performance
Figure 2 shows the performance of humans and LLMs in the Defaults condition as a function of their performance on MMLU [^4] MMLU scores are few-shot for GPT-4 and 5-shot for other models. The reported human baseline is the estimate for human experts given by Hendrycks et al. [31]. The score for Pythia 12B could not be found and so we use the reported value for Pythia 6.9B Tulu., a widely-used language competency benchmark. Increasing MMLU score is associated with more matches to reference answers on the Defaults condition. This steadily increasing performance is presumed to correlate with the scale of model parameters and training data [32]. Smaller models do not perform competitively (Pythia-12B obtains a reference match of 0.0, Falcon 40B 0.1, GPT-3 0.5, and Claude 2 0.6). We thus focus our remaining analysis on comparing human subjects to GPT-4, Claude 3, and Llama-405B. In the Defaults condition, none of GPT-4 (coef=-0.7696, z=-1.659, p=0.097), Claude 3 (coef=-0.6131, z=-1.299, p=0.194), or Llama-405B (coef=-0.7696, z=-1.659, p=0.097) performs significantly worse than human subjects. These top three models are notably similar in both their MMLU score and their corresponding performance in the Defaults condition of our experiment.
@@
Figure 2: Human and LLM reference match proportion in the Defaults condition, relative to performance on the MMLU benchmark. Models in blue are not instruction-tuned while models in peach are. Error bars show standard errors. MMLU scores are collated from various sources.
4.1.2 Sensitivity to Question Presentation
We next compare performance in Defaults to performance in our Permuted Questions, Permuted Pairs, and Distracted settings. Recall that these conditions are designed to probe sensitivity to the order and presentation of the analogy. If humans and models exhibit different changes in performance relative to Defaults in these conditions, this would be indicative of differences in the mechanisms being used to infer the analogy from the provided text.
@@ 4.1.2 Sensitivity to Question Presentation
Figure 3 compares humans to high-performing LLMs in the Defaults and Permuted Pairs conditions. LLM performance drops in the Permuted Pairs condition, while humans seem equally able to infer the mapping regardless of word presentation order. This effect is significant for Claude 3 (coef=-1.7802, z=-4.217, p<0.001), GPT-4 (coef=-1.6796, z=-3.975, p<0.001), and Llama-405B (coef=-2.3381, z=-5.500, p<0.001). The remaining control conditions and data for all tested models are shown in Figure 15 of the Appendix. In these conditions, we find that humans and models are roughly equally affected. For example, the proportion of matches with the reference answers in the Distracted condition drops by approximately 0.25 for all four subject types.
@@
Figure 3: Human and LLM reference match proportion in the Defaults and Permuted Pairs conditions. Error bars show standard errors.
@@ 4.1.3 Effect of Semantic Structure on Reasoning
We next investigate more directly the extent to which humans and LLMs leverage semantic structure in order to complete our analogy tasks by comparing performance in Defaults to performance in the Only RHS, Randoms, and Random Finals conditions. Recall (from §2.1) that we interpret high performance in the Only RHS condition as evidence that a subject is able to complete the questions based only on the pattern in the target domain. We take the performance difference between the Only RHS condition and either the Randoms or Random Finals condition to be a measure of the subject’s bias toward using the semantic structure: a large performance degradation indicates that the subject preferred to consider the LHS as meaningful even when ignoring the LHS entirely would have arguably afforded a simpler strategy for producing responses.
@@
Both humans and some models are readily able to infer the reference answer in the Only RHS condition, as shown in Figure 4. Match to reference answers is approximately 0.8 for Claude 3 with human subjects and GPT-4 slightly higher at 0.9. Llama-405B achieves lower but still reasonable matches to the reference of approximately 0.5. GPT-4 is not significantly different from humans in this condition (coef=0.1178, z=0.223, p=0.824), while Claude 3 (coef=-0.9130, z=-1.994, p=0.046) and Llama-405B (coef=-2.0294, z=-4.648, p<0.001) are significantly worse than human subjects. Thus, both humans and some LLMs are able to complete the task without the guidance of the left hand side.
Considering this, we next look at the performance degradation associated with encountering incoherent semantic structure on the left hand side. Humans exhibit a modest decrease in matches to the reference answers of about 0.15 in the Randoms and Random Permuted Pairs conditions relative to Defaults. Claude-3 and GPT-4, however, exhibit much larger drops: Claude 3 decreases by approximately 0.5 relative to Defaults, while GPT-4 decreases by 0.6 and 0.4 in the Randoms and Random Permuted Pairs conditions. The performance of Llama-405B is comparably low in these two conditions, although reducing from a lower score in the Only RHS condition. Across these two conditions, GPT-4 (coef=-2.1972, z=-5.211, p<0.001), Claude 3 (coef=-2.0680, z=-4.960, p<0.001), and Llama-405B (coef=-2.7383, z=-6.078, p<0.001) perform significantly worse than humans.
@@
Figure 4: Human and LLM reference match proportion in Only RHS, Randoms, and Random Finals conditions. Data from the Random Permuted Pairs condition is shown in Figure 15 of the Appendix. Error bars show standard errors.
@@
From this we conclude that human subjects are able to easily identify when the left hand side contains no useful semantic structure to leverage. In such cases, they are able to employ a strategy that only relies on the right hand side. By contrast, models do not seem capable of easily identifying the lack of informativeness of the left hand side in these conditions, as they do not use the strategy of only attending to the right hand side, even though they show their capability of using this strategy when no left hand side is present. This suggests mechanistic differences between how human subjects and models process this task.
@@
Although the performance of human subjects does not drop notably in the Randoms condition compared to the Only RHS condition, it does drop by a wide margin in the Random Finals condition. In this condition, correspondence with reference answers is approximately 0.5 lower than in the Only RHS condition. This further suggests that the semantic relatedness of the left hand side affects the strategy of human subjects: when the left hand side is clearly unrelated, the information it provides is discarded, but when much of the left hand side appears related, the information is not discarded and the random final word of the source domain prompts an incorrect answer from human subjects. Models also show a large drop in performance in the Random Finals condition relative to Only RHS, with Claude 3 dropping by 0.5, GPT-4 dropping by 0.8, and Llama-405B dropping by 0.4 (with this smaller drop being due to a lower starting point of 0.5). Simple effects analysis shows that Claude 3 (coef=-1.0464, z=-2.799, p=0.005), GPT-4 (coef=-2.7850, z=-5.168, p< 0.001), and Llama-405B (coef=-2.0229, z=-4.625, p< 0.001) are significantly worse than humans in the Random Finals condition. However, we see this difference as less informative given that all models drop in performance across all the random conditions relative to their own performance in the Only RHS condition.
4.1.4 Diagnosing the Use of an RHS-Only Heuristic
To clarify whether subjects actually make use of left-right relations or only complete right-side patterns in the Semantic Structure experiment, we use our $2\times n$ variant of the Defaults condition, preventing the use of only right-hand terms.
Results are shown in Figure 5. Human subjects and Claude 3 exhibit similar performance, with accuracies of approximately 0.7. GPT-4 and Llama-405B, however, both attain a much lower reference match ratio close to 0.35. Simple effects analysis shows that GPT-4 (coef=-1.3669, z=-3.065, p=0.002) and Llama-405B (coef=-1.2550, z=-2.843, p=0.004) obtain significantly worse match to the reference answers than human subjects, while the reference match of Claude 3 does not differ significantly from human subjects (coef=0.2111, z=0.467, p=0.640).
Figure 5: Human and LLM reference match proportion in the $2\times n$ condition followup, with Defaults condition performance for reference. Error bars show standard errors.